OpenStudy (anonymous):

1/2 divided by 1/2 what i did was i 1/2 x 2/1= 1 but if u divided it i comes to be 1/2 divided 1/2 = 0.25 which one is right ?

OpenStudy (anonymous):

@R-Y-P

OpenStudy (anonymous):

Ya she's correct

OpenStudy (triciaal):

anything divided by itself is one 2/2 =1 4/4 =1 4.5/4.5 = 1 1/2 divided by 1/2 = 1